# Uniform Hyperbolicity and Symbolic Dynamics: Markov Partitions, Shadowing, and the Coding of Axiom A Systems

### 19 Apr 2026 v1

@@ 4.6. Hölder Dependence on Base Point
The stable manifolds depend Hölder continuously on the base point. This regularity is what allows the coding map constructed in Section 9 to be Hölder, which is essential for transferring the spectral theory of Part I to the smooth setting.
@@ 7.3. Anosov Closing Lemma
Let $y\in M$ be arbitrary and $\gamma>0$. Let $\beta=\gamma/2$ and let $\alpha>0$ be the constant from the Anosov Closing Lemma (Theorem 7.5) corresponding to this $\beta$. Since $M=\Omega(f)$ (every point is nonwandering for an Anosov diffeomorphism, as stable and unstable manifolds through any point are dense in $M$), the point $y$ is nonwandering. By definition, there exists $n>0$ such that $f^{n}(B(y,\alpha/2))\cap B(y,\alpha/2)\neq\emptyset$. Choose $x\in B(y,\alpha/2)$ with $f^{n}(x)\in B(y,\alpha/2)$. Then $d(f^{n}(x),x)\leq d(f^{n}(x),y)+d(y,x)<\alpha/2+\alpha/2=\alpha$. By Theorem 7.5, there exists a periodic point $p\in M$ of period $n$ with $d(p,x)\leq\beta=\gamma/2$. Thus $d(p,y)\leq d(p,x)+d(x,y)<\gamma/2+\alpha/2\leq\gamma$ for $\alpha\leq\gamma$. ∎
@@ 7.4. Specification Property
For periodic specification: close the pseudo-orbit by connecting the last segment back to the first using the mixing property, with total period $p=a_{k+1}$ where $a_{k+1}-b_{k}\geq N$. The resulting bi-infinite periodic pseudo-orbit is $\alpha$-shadowed by some $x\in\Omega$. By the Anosov Closing Lemma (Theorem 7.5), since $d(f^{p}(x),x)\leq 2\beta<\varepsilon_{\exp}$, there is a periodic point $q$ of period $p$ with $d(f^{i}(q),f^{i}(x))\leq\varepsilon/2$ for all $i$. Then $d(f^{a_{j}+i}(q),f^{i}(x_{j}))\leq d(f^{a_{j}+i}(q),f^{a_{j}+i}(x))+d(f^{a_{j}+i}(x),f^{i}(x_{j}))\leq\varepsilon/2+\varepsilon/2=\varepsilon$. ∎
@@ 7.5. Quantitative Periodic Orbit Approximation
This subsection quantifies how many periodic orbits of each period exist and how well they approximate invariant measures. These estimates underlie the counting and equidistribution statements needed in Parts IV–VI.
@@ 9. Symbolic Dynamics
This section establishes the symbolic coding of Axiom A diffeomorphisms via Markov partitions, providing the essential bridge between smooth dynamics and the symbolic dynamics developed in Parts I and II Thiam (2025a, b). We prove quantitative bounds on the coding map and characterize the exceptional set where the coding fails to be injective.
@@ 9.3. Quantitative Coding Estimates
This subsection provides the quantitative bounds on the coding map that are used in Parts IV–VI: Hölder continuity with explicit exponent, and control on the exceptional set where the map fails to be injective.
@@ 9.5. Markov Property Extension
The Markov property is initially established on a dense subset of $\Omega$. This subsection extends it to all points, which is needed for the entropy calculation and for the unique ergodicity statements in Part IV.
@@ 11. Conclusion
Part IV of this series uses this coding to develop the transfer operator theory on manifolds, construct SRB measures, establish the Gibbs Equivalence Theorem, and prove structural stability with quantitative Hölder bounds on conjugacies. Parts V and VI develop the statistical limit theorems and the structural consequences for higher-dimensional systems.
@@ A.3. Spectral Theory for Quasi-compact Operators
We recall the spectral decomposition for quasi-compact bounded operators on a Banach space. This is used in Parts IV–V, where the Ruelle transfer operator on $C^{\alpha}$-manifolds is shown to be quasi-compact.
@@ A.5. Geometric Measure Theory
We state the definition of Hausdorff dimension and recall Bowen’s equation, which links the Hausdorff dimension of a conformal repeller to the zero of an associated pressure function. These tools are used in the dimension-theoretic consequences developed in Parts V–VI.

### 23 Apr 2026 v2

@@ 4.6. Hölder Dependence on Base Point
The stable manifolds depend Hölder continuously on the base point. This regularity is what allows the coding map constructed in Section 9 to be Hölder, which is essential for transferring the spectral theory of Part I Thiam (2026a) to the smooth setting.
@@ 7.3. Anosov Closing Lemma
Let $y\in M$ be arbitrary and $\gamma>0$. Let $\beta=\gamma/2$ and let $\alpha>0$ be the constant from the Anosov Closing Lemma (Proposition 7.5) corresponding to this $\beta$. Since $M=\Omega(f)$ (every point is nonwandering for an Anosov diffeomorphism, as stable and unstable manifolds through any point are dense in $M$), the point $y$ is nonwandering. By definition, there exists $n>0$ such that $f^{n}(B(y,\alpha/2))\cap B(y,\alpha/2)\neq\emptyset$. Choose $x\in B(y,\alpha/2)$ with $f^{n}(x)\in B(y,\alpha/2)$. Then $d(f^{n}(x),x)\leq d(f^{n}(x),y)+d(y,x)<\alpha/2+\alpha/2=\alpha$. By Proposition 7.5, there exists a periodic point $p\in M$ of period $n$ with $d(p,x)\leq\beta=\gamma/2$. Thus $d(p,y)\leq d(p,x)+d(x,y)<\gamma/2+\alpha/2\leq\gamma$ for $\alpha\leq\gamma$. ∎
@@ 7.4. Specification Property
For periodic specification: close the pseudo-orbit by connecting the last segment back to the first using the mixing property, with total period $p=a_{k+1}$ where $a_{k+1}-b_{k}\geq N$. The resulting bi-infinite periodic pseudo-orbit is $\alpha$-shadowed by some $x\in\Omega$. By the Anosov Closing Lemma (Proposition 7.5), since $d(f^{p}(x),x)\leq 2\beta<\varepsilon_{\exp}$, there is a periodic point $q$ of period $p$ with $d(f^{i}(q),f^{i}(x))\leq\varepsilon/2$ for all $i$. Then $d(f^{a_{j}+i}(q),f^{i}(x_{j}))\leq d(f^{a_{j}+i}(q),f^{a_{j}+i}(x))+d(f^{a_{j}+i}(x),f^{i}(x_{j}))\leq\varepsilon/2+\varepsilon/2=\varepsilon$. ∎
@@ 7.5. Quantitative Periodic Orbit Approximation
This subsection quantifies how many periodic orbits of each period exist and how well they approximate invariant measures. These estimates underlie the counting and equidistribution statements needed in Parts IV–VI Thiam (2026d, e, f).
@@ 9. Symbolic Dynamics
This section establishes the symbolic coding of Axiom A diffeomorphisms via Markov partitions, providing the essential bridge between smooth dynamics and the symbolic dynamics developed in Parts I and II Thiam (2026a, b). We prove quantitative bounds on the coding map and characterize the exceptional set where the coding fails to be injective.
@@ 9.3. Quantitative Coding Estimates
This subsection provides the quantitative bounds on the coding map that are used in Parts IV–VI Thiam (2026d, e, f): Hölder continuity with explicit exponent, and control on the exceptional set where the map fails to be injective.
@@ 9.5. Markov Property Extension
The Markov property is initially established on a dense subset of $\Omega$. This subsection extends it to all points, which is needed for the entropy calculation and for the unique ergodicity statements in Part IV Thiam (2026d).
@@ 11. Conclusion
Part IV Thiam (2026d) of this series uses this coding to develop the transfer operator theory on manifolds, construct SRB measures, establish the Gibbs Equivalence Theorem, and prove structural stability with quantitative Hölder bounds on conjugacies. Parts V and VI Thiam (2026e, f) develop the statistical limit theorems and the structural consequences for higher-dimensional systems.
@@ A.3. Spectral Theory for Quasi-compact Operators
We recall the spectral decomposition for quasi-compact bounded operators on a Banach space. This is used in Parts IV–V Thiam (2026d, e), where the Ruelle transfer operator on $C^{\alpha}$-manifolds is shown to be quasi-compact.
@@ A.5. Geometric Measure Theory
We state the definition of Hausdorff dimension and recall Bowen’s equation, which links the Hausdorff dimension of a conformal repeller to the zero of an associated pressure function. These tools are used in the dimension-theoretic consequences developed in Parts V–VI Thiam (2026e, f).